# MME: A Comprehensive Evaluation Benchmark for Multimodal Large Language Models

### v1

@@ 2.3.1 Perception Tasks
We argue that perception is one of the most fundamental capabilities of MLLM, and the lack of perception will easily lead to the object hallucination problem Li et al. (2023d). That is, MLLM will answer questions based on its own fantasies rather than based on the realistic content of the image, as displayed in Fig. 4.
Coarse-Grained Recognition. The contents of coarse-grained recognition include the existence of common objects, and their count, color, and position. The images are sampled from COCO Lin et al. (2014), but the instruction-answer pairs are all manually constructed, rather than directly using publicly available annotations. Even if MLLMs have seen these COCO images, our manually prepared pairs are not presented in their training sets. This requires MLLM to be able to understand the instructions and infer corresponding answers. In each perception subtask of existence, count, color, and position, we prepare 30 images with 60 instruction-answer pairs.
Fine-Grained Recognition. The fine-grained recognition is more about testing the knowledge resources of MLLM. The subtasks consist of recognizing movie posters, celebrities, scenes, landmarks, and artworks, containing 147, 170, 200, 200, and 200 images respectively. For the celebrities, we plot a red box to a person with a clearly visible face in the image, and the corresponding instruction is “Is the actor inside the red box named [celebrity name]? Please answer yes or no.” Similar with the above coarse-grained recognition, the images of these subtasks are from publicly available datasets Huang et al. (2020); Mao et al. (2017, 2019); Zhou et al. (2014); Weyand et al. (2020) and all of the instructions are manually designed.
OCR. Optical Character Recognition (OCR) is also a foundational capability of MLLM, serving for subsequent text-based tasks such as text translation and text understanding. The images are from Liu et al. (2019) and the instruction-answer pairs are manually designed. Considering that MLLM is still in its infancy, we only choose the relatively simple samples in this version of MME. The numbers of image and instruction-answer pairs are 20 and 40, respectively.
@@ 2.3.2 Cognition Tasks
We evaluate if any MLLM can carry out further logical reasoning after perceiving the image, which is the most fascinating aspect of MLLM over previous traditional methods. In order to infer the correct answer, MLLM needs to follow the instruction, perceive the contents of the image, and invoke the knowledge reserved in LLM, which is much more challenging than the single perception tasks. Examples of the following subtasks are shown in Fig. 1.
Commonsense Reasoning. Unlike the ScienceQA dataset Lu et al. (2022) that requires specialized knowledge, the commonsense refers to the basic knowledge in daily life. For example, given a photo of a down jacket, asking MLLM whether it is appropriate to wear the cloth when it is cold (or hot). These are basic knowledge that humans can judge instantly without complex step-by-step reasoning. Therefore, we expect MLLM to perform well in a zero-short setting. The images are all manually photographed or generated by diffusion models, and the instruction-answer pairs are all manually designed. There are a total of 70 images and 140 instruction-answer pairs.
Numerical Calculation. It requires MLLM to be able to read the arithmetic problem in the image and output the answer in an end to end way, which has been demonstrated in Huang et al. (2023). In this version, we only consider relatively easy arithmetic problems, such as addition, subtraction, multiplication, and division. There are a total of 20 images and 40 instruction-answer pairs. The images are all manually taken, and the instruction-answer pairs are all manually designed.
Text Translation. Considering that the MLLM git (2023i) supports both English and Chinese, we set the text translation task. It requires MLLM to translate the Chinese written in an image to the corresponding English. In this version, we only design basic translation problems, which will be updated according to the development of MLLM in the future. The images of this part are all manually taken, and the instruction-answer pairs are all manually designed. There are a total of 20 images and 40 instruction-answer pairs.
Code Reasoning. It requires MLLM to read the code in the images and automatically complete logical operation inside the code. A similar task that writes website code based on an image has been demonstrated in Zhu et al. (2023). The images are all manually taken, and the instruction-answer pairs are all manually designed. We only set basic code problems in this version. There are in total 20 images and 40 instruction-answer pairs.
@@ 3 Experiments
Model Existence Count Position Color OCR ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ BLIP-2 86.67 73.33 75.00 60.00 56.67 16.67 81.67 66.67 70.00 40.00 LLaVA 50.00 0.00 50.00 0.00 50.00 0.00 51.67 3.33 50.00 0.00 MiniGPT-4 75.00 60.00 66.67 56.67 56.67 33.33 71.67 53.33 62.50 35.00 mPLUG-Owl 73.33 46.67 50.00 0.00 50.00 0.00 51.67 3.33 55.00 10.00 LLaMA-AdapterV2 76.67 56.67 58.33 6.67 43.33 3.33 55.00 16.67 57.50 15.00 InstructBLIP 95.00 90.00 80.00 63.33 53.33 13.33 83.33 70.00 57.50 15.00 VisualGLM-6B 61.67 23.33 50.00 0.00 48.33 0.00 51.67 3.33 42.50 0.00 Otter 53.33 6.67 50.00 0.00 50.00 0.00 51.67 3.33 50.00 0.00 Multimodal-GPT 46.67 10.00 51.67 6.67 45.00 13.33 55.00 13.33 57.50 25.00 PandaGPT 56.67 13.33 50.00 0.00 50.00 0.00 50.00 0.00 50.00 0.00
Table 1: Evaluation of coarse-grained recognition and OCR.
Model Poster Celebrity Scene Landmark Artwork ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ BLIP-2 79.25 62.59 68.53 37.06 81.25 64.00 79.00 59.00 76.50 60.00 LLaVA 50.00 0.00 48.82 0.00 50.00 0.00 50.00 0.00 49.00 0.00 MiniGPT-4 49.32 19.73 58.82 24.71 68.25 45.50 59.75 30.50 56.25 27.00 mPLUG-Owl 77.89 57.14 66.18 34.12 78.00 57.50 86.25 73.00 63.25 33.00 LLaMA-AdapterV2 52.72 10.88 55.00 21.18 68.75 44.50 53.00 9.00 52.50 14.50 InstructBLIP 74.15 49.66 67.06 34.12 84.00 69.00 59.75 20.00 76.75 57.50 VisualGLM-6B 54.42 12.24 50.88 2.35 81.75 64.50 59.75 24.00 55.75 20.00 Otter 45.24 0.00 50.00 0.00 55.00 14.50 52.00 4.50 48.00 5.50 Multimodal-GPT 45.24 17.01 49.12 24.12 50.50 17.50 50.50 23.00 46.00 12.00 PandaGPT 56.80 19.73 46.47 10.59 72.50 45.50 56.25 13.50 50.25 1.00
Table 2: Evaluation of fine-grained recognition.
Model Commonsense Reasoning Numerical Calculation Text Translation Code Reasoning ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ BLIP-2 68.57 41.43 40.00 0.00 55.00 10.00 55.00 20.00 LLaVA 49.29 11.43 50.00 0.00 52.50 5.00 50.00 0.00 MiniGPT-4 58.57 34.29 47.50 20.00 42.50 15.00 67.50 45.00 mPLUG-Owl 59.29 24.29 50.0 10.00 60.00 20.00 47.50 10.00 LLaMA-Adapter V2 54.29 14.29 52.50 5.00 52.50 5.00 52.50 10.00 InstructBLIP 75.00 54.29 35.00 5.00 55.00 10.00 47.50 0.00 VisualGLM-6B 45.71 12.86 45.00 0.00 55.00 10.00 50.00 0.00 Otter 48.57 10.00 47.50 10.00 55.00 10.00 50.00 0.00 Multimodal-GPT 45.71 5.71 50.00 20.00 50.00 5.00 45.00 10.00 PandaGPT 56.43 17.14 50.00 0.00 52.50 5.00 47.50 0.00
Table 3: Evaluation of cognition.
In this section, we conduct massive experiments on our MME benchmark to evaluate a total of 10 open-source MLLMs, including BLIP-2, LLaVA, MiniGPT-4, mPLUG-Owl, LLaMA-Adapter-v2, Otter, Multimodal-GPT, InstructBLIP, VisualGLM-6B, and PandaGPT. Except BLIP-2, other models have been fine-tuned on their instruction tuning datasets.
3.1 Models
All experiments are conducted on NVIDIA V100 GPUs, and we use as large models as possible. BLIP-2 Li et al. (2023c) focuses on the basic pre-training, where the image encoder and the LLM are all frozen and a light-weight Q-Former is trained for multimodal representation alignment and vision-to-language generation. Despite not performing the multimodal instruction tuning, BLIP-2 is still able to follow instructions with the LLM’s own capabilities. The version under investigation is “blip2-pretrain-flant5xxl” git (2023a).
LLaVA Liu et al. (2023) is a pioneer in bringing multimodal instruction tuning to MLLM. The tuning data are generated by language-only GPT-4 through carefully designed prompts. The instructions consist of three types, including conversation, detailed description, and complex reasoning. A projection layer and LLM are updated during training. The tested version is “LLaVA-7B-v0” git (2023d).
MiniGPT-4 Zhu et al. (2023) performs multimodal instruction tuning on the pre-trained BLIP-2, where only a linear layer is updated. The instructions are mainly based on the image caption task, such as “Describe this image in detail.” We use “minigpt4-aligned-with-vicuna13b” git (2023e) for testing.
mPLUG-Owl Ye et al. (2023) uses both language-only instruction data and the multimodal instruction data from LLaVA. The visual encoder, the proposed visual abstractor, and LLM (insert a LoRA module Hu et al. (2021)) are all updated during training. We test “mplug-owl-llama-7b” git (2023f).
LLaMA-Adapter V2 Gao et al. (2023) is only trained on language-only instruction data and image-text pairs via a parameter-efficient manner. Additional expert models, such as detection, can be integrated to boost reasoning. The version of “LLaMA-7B” git (2023c) is used in testing.
InstructBLIP Dai et al. (2023) is based on BLIP-2 and reorganizes 26 existing public datasets like image caption and VQA as instruction tuning format. Only the Q-Former is updated during training. “blip2-instruct-flant5xxl” git (2023b) is tested.
VisualGLM-6B git (2023i) is an open-source MLLM but with unclear training details by now. It supports both English and Chinese. We test “VisualGLM-6B”  git (2023i).
Otter Li et al. (2023b) mixes multimodal in-context learning and multimodal instruction tuning. The framework is based OpenFlamingo Awadalla et al. (2023), whose perceiver module and parts of LLM are updated during training. The instruction data are from VQAv2 Antol et al. (2015), GQA Hudson and Manning (2019), LLaVA, and a not public video dataset. The version of “OTTER-9B-LA-InContext” git (2023g) is tested.
Multimodal-GPT Gong et al. (2023) is also based on OpenFlamingo and uses both language-only and multimodal instruction data. The latter includes data from LLaVA, Mini-GPT4, and various image caption and VQA datasets. Only the LoRA module in LLM is updated during training. The tested version is “Multimodal-GPT-9B”.
PandaGPT Su et al. (2023) uses multimodal instruction tuning data from LLaVA and MiniGPT-4, and only a linear projection layer as well as the LoRA module in LLM are trainable. The pretrained ImageBind Girdhar et al. (2023) is adopted as the multimodal encoder, which has an attribute of modality alignment. “pandagpt-7b-max-len-512” git (2023h) is used for testing.
@@ 3.2.1 Perception
There are a total of 10 subtasks for the evaluation of the perception ability, from the perspectives of coarse-grained recognition, fine-grained recognition, and OCR. Figs. 2 (3)-(6) show the score leaderboards of individual coarse-grained recognition subtasks. With respect to the object existence, InstructBLIP gets the highest score 185, with a 95% accuracy and a 90% accuracy+ listed in Table 1. Contrastively, the second place BLIP-2 and the third place MiniGPT-4 lag behind InstructBLIP by 25 and 50 scores, respectively. For the object count, position, and color, InstructBLIP, BLIP-2, and MiniGPT-4 still make the top three, which is reasonable because the object existence is the basis of object attributes. What is interesting is that both InstructBLIP and MiniGPT-4 are instruction finetuned on the pre-trained BLIP-2. Note that in the four coarse-grained subtasks, these MLLMs get the worst results on object position, indicating that the current models are not sensitive enough to the position information.
Figs. 2 (7)-(11) show the score leaderboards of individual fine-grained recognition subtasks. Regarding to poster recognition, BLIP-2, mPLUG-Owl, and InstructBLIP are the top three. It is interesting that mPLUG-Owl underperforms in the coarse-grained recognition, but now it exhibits good. This implies that our division of coarse-grained and fine-grained is reasonable and enables us to examine different aspects of MLLMs. For the celebrity recognition, BLIP-2, InstructBLIP, and mPLUG-Owl still take the top three with similar scores. For the scene recognition, InstructBLIP, VisualGLM-6B, and BLIP-2 ahead of other MLLMs. This is the first time VisualGLM-6B has broken into the top three. Also noteworthy is PandaGPT, which scores more than a hundred for the first time. For the landmark recognition, top three places are taken by mPLUG-Owl, BLIP-2, and MiniGPT-4 respectively, of which mPLUG-Owl gets the top spot. For the artwork recognition, BLIP-2, InstructBLIP, and mPLUG-Owl exceed other counterparts, where the first two scores are similar, leading the third by 38+. With respect to OCR listed in Fig. 2 (12), BLIP-2, MiniGPT-4, and Multimodal-GPT get the top three with scores of 110, 97.5, and 82.5 respectively, in which the first two perform more outstanding.
Figure 4: Common problems revealed in our experiments. [Y]/[N] means the ground truth answer is yes/no. [R] is the generated answer of MLLM.
As can be seen in Fig. 2 (1), in the leaderboard of the whole perception recognition, BLIP-2, InstructBLIP, and MiniGPT-4 come in top three, closely followed by mPLUG-Owl. Their perception scores are 1293.84, 1212.82, 1010.66, and 966.32, respectively. Among them, mPLUG-Owl shows a clear bias, good at fine-grained recognition, but not good at coarse-grained recognition. On the contrary, MiniGPT-4 has worse fine-grained recognition than coarse-grained recognition, while the discrepancy is not that significant as mPLUG-Owl. Comparatively, the overall performances of BLIP-2 and InstructBLIP are more stable.
@@ 3.2.2 Cognition
There are four subtasks for the evaluation of the cognition ability, including commonsense reasoning, numerical calculation, text translation, and code reasoning. Figs. 2 (13)-(16) plot the score leaderboards of individual subtasks. In terms of the commonsense reasoning, the “ever-victorious generals” InstructBLIP, BLIP-2, and MiniGPT-4 exceed other MLLMs, especially InstructBLIP, which gets a score of 129.29. With respect to numerical calculation and text translation, although the designed questions are moderate in difficulty, these MLLMs do not perform satisfactorily, with no score beyond 80. This suggests that MLLMs have a lot of room for improvement in these capabilities. For the code reasoning, MiniGPT-4 achieves a high score of 112.5, far ahead of other counterparts. This is reminiscent of the example it showed of writing website codes from an image. For all of the cognition tasks, MiniGPT-4, InstructBLIP, and BLIP-2 win the gold, silver, and bronze medals respectively, as shown in Fig. 2 (2).
@@ 4 Analysis
The third problem is a lack of reasoning. In the third row of Fig. 4, we can see from the red text that the MLLM already knows that the first image is not an office place, but still gives an incorrect answer of “yes”. Analogously, in the second image, the MLLM has calculated the right arithmetic result, but finally delivers a wrong answer. These phenomena indicate that the logic chain is broken during the reasoning process of MLLMs. Adding CoT prompts, such as “Let’s think step by step” Driess et al. (2023), may yield better results. We look forward to a further in-depth research.
The fourth problem is object hallucination following instructions, which is exemplified in the fourth row of Fig. 4. When the instruction contains descriptions of an object that does not appear in the image, the MLLM will imagine that the object exists and ultimately gives a “yes” answer. Such a case of constantly answering “yes” results in an accuracy about 50% and an accuracy+ about 0, as shown in Tables 1, 2, and 3. This suggests an urgent need to suppress hallucinations, and the community should take into account of the reliability of the generated answers.
@@ 5 Conclusion
This paper has presented the first MLLM evaluation benchmark MME that has four distinct characteristics in terms of task type, data source, instruction design, quantitative statistics. We evaluate 10 advanced MLLMs on MME and the experimental results show that there is still a large room to improve. We also summarize the common problem raised in experimental results, providing valuable guidance for the development of MLLM.

### v4

@@ 2.3.1 Perception Tasks
We argue that perception is one of the most fundamental capabilities of MLLMs, and the lack of perception will easily lead to the object hallucination problem [50, 25]. That is, MLLM will answer questions based on its own fantasies rather than based on the realistic content of the image, as displayed in Fig. 4.
Coarse-Grained Recognition. The contents of coarse-grained recognition include the existence of common objects, and their count, color, and position. The images are sampled from COCO [26], but the instruction-answer pairs are all manually constructed, rather than directly using publicly available annotations. Even if MLLMs have seen these COCO images, our manually prepared pairs are not presented in their training sets. This requires MLLMs to be able to understand the instructions and infer corresponding answers. In each perception subtask of existence, count, color, and position, we prepare 30 images with 60 instruction-answer pairs.
Fine-Grained Recognition. The fine-grained recognition is more about testing the knowledge resources of MLLMs. The subtasks consist of recognizing movie posters, celebrities, scenes, landmarks, and artworks, containing 147, 170, 200, 200, and 200 images respectively. For the celebrities, we plot a red box to a person with a clearly visible face in the image, and the corresponding instruction is “Is the actor inside the red box named [celebrity name]? Please answer yes or no.” Similar with the above coarse-grained recognition, the images of these subtasks are from publicly available datasets [19, 34, 35, 58, 44] and all of the instructions are manually designed.
OCR. Optical Character Recognition (OCR) is also a foundational capability of MLLMs, serving for subsequent text-based tasks such as text translation and text understanding. The images are sampled from [30] and all of the instruction-answer pairs are manually designed. Considering that MLLMs are still in its infancy, we only choose the relatively simple samples in this version of MME. The numbers of image and instruction-answer pairs are 20 and 40, respectively.
@@ 2.3.2 Cognition Tasks
We evaluate if any MLLM can carry out further logical reasoning after perceiving the image, which is the most fascinating aspect of MLLMs over previous traditional methods. In order to infer the correct answer, MLLMs need to follow the instruction, perceive the contents of the image, and invoke the knowledge reserved in LLMs, which is much more challenging than the single perception tasks. Examples of the following subtasks are shown in Fig. 1.
Commonsense Reasoning. Unlike the ScienceQA dataset [32] that requires specialized knowledge, the commonsense refers to the basic knowledge in daily life. For example, given a photo of a down jacket, asking MLLMs whether it is appropriate to wear the cloth when it is cold (or hot). These are basic knowledge that humans can judge instantly without complex step-by-step reasoning. Therefore, we expect MLLMs to perform well in a zero-short setting. The images are all manually photographed or generated by diffusion models, and the instruction-answer pairs are all manually designed. There are a total of 70 images and 140 instruction-answer pairs.
Numerical Calculation. It requires MLLMs to be able to read the arithmetic problem in the image and output the answer in an end to end way, which has been demonstrated in [20]. In this version, we only consider relatively easy arithmetic problems, such as addition and multiplication. There are 20 images and 40 instruction-answer pairs. The images are all manually taken, and the instruction-answer pairs are all manually designed.
Text Translation. Considering that the MLLM [5] supports both English and Chinese, we set the text translation subtask. It requires MLLMs to translate the Chinese written in an image to the corresponding English. In this version, we only design basic translation problems, which will be updated according to the development of MLLMs in the future. The images of this part are all manually taken, and the instruction-answer pairs are all manually designed. There are a total of 20 images and 40 instruction-answer pairs.
Code Reasoning. It requires MLLMs to read the code in the images and automatically complete logical operation inside the code. A similar task that writes website code based on an image has been demonstrated in [59]. The images are all manually taken, and the instruction-answer pairs are all manually designed. We only set basic code problems in this version. There are in total 20 images and 40 instruction-answer pairs.
@@ 3 Experiments
In this section, a total of 30 MLLMs are evaluated on our MME benchmark, including BLIP-2 [23], InstructBLIP [12], MiniGPT-4 [59], PandaGPT [39], Multimodal-GPT [15], VisualGLM-6B [5], ImageBind-LLM [17], VPGTrans [53], LaVIN [33], mPLUG-Owl [48], Octopus [3], Muffin [51], Otter [22], LRV-Instruction [28], Cheetor [24], LLaMA-Adapter-v2 [14], GIT2 [41], BLIVA [18], Lynx [52], MMICL [54], GPT-4V [37], Skywork-MM [4], mPLUG-Owl2 [48], Qwen-VL-Chat [9], XComposer-VL [7], LLaVA [29], Lion [2], SPHINX [27], InfMLLM [1], and WeMM [6].
@@ 3.1.1 Perception
There are a total of 10 subtasks for the evaluation of the perception ability, from the perspectives of coarse-grained recognition, fine-grained recognition, and OCR. Figs. 2 (3)-(6) show the score leaderboards of individual coarse-grained recognition subtasks. With respect to the object existence, Otter, Lynx, WeMM, Muffin, and SPHINX get the highest score 195, with a 98.33% accuracy and a 96.67% accuracy+ listed in Table 1. Contrastively, the second place, including GIT2, XComposer-VL, Lion, GPT-4V, and etc, lag behind the first place only by 5 scores. The results show that these models already have a good performance on object existence. For the object count, position, and color, Muffin, Lion (parallel with SPHINX), and InfMLLM make the top one, respectively. It suggests that different models have their own strengths. Note that in the four coarse-grained subtasks, these MLLMs get the worst results on object position, indicating that the current models are not sensitive enough to the position information.
Figs. 2 (7)-(11) display the score leaderboards of individual fine-grained recognition subtasks. Regarding to poster recognition, GPT-4V, Lion, and Qwen-VL-Chat are the top three. It is interesting that Qwen-VL-Chat relatively underperforms in the coarse-grained recognition, but now it exhibits good. This implies that our division of coarse-grained and fine-grained is reasonable, enabling us to examine different aspects of MLLMs. For the celebrity recognition, WeMM, SPHINX, and Otter take the top three with similar scores. It is worth noting that GPT-4V refuses to answer questions that involve individuals, resulting in a zero score in the celebrity subtask. For the scene recognition, WeMM, InfMLLM, and Lynx ahead of other MLLMs. This is the first time InfMLLM and Lynx have broken into the top three in the fine-grained recognition subtasks. For the landmark recognition, top three places are taken by Lion, WeMM, and LLaVA respectively, of which Lion gets the top spot. For the artwork recognition, WeMM, GPT-4V, and GIT2 exceed other counterparts, where the last two scores are similar. Note that GPT-4V declines to answer some questions about private art collection, which lowers its score. With respect to OCR listed in Fig. 2 (12), GPT-4V, Skywork-MM, and WeMM get the top three with scores of 185, 162.5, and 147.5 respectively. GPT-4V presents a huge advantage, leading the other two models by 22+ socres. As presented in Fig. 2 (1), in the leaderboard of the whole perception recognition, WeMM, InfMLLM, and SPHINX come in top three, closely followed by Lion, LLaVA, and XComposer-VL.
Figure 4: Common problems revealed in experiments. [Y]/[N] means the ground truth answer is yes/no. [R] is the generated answer.
@@ 3.1.2 Cognition
There are four subtasks for the evaluation of the cognition ability, including commonsense reasoning, numerical calculation, text translation, and code reasoning. Figs. 2 (13)-(16) plot the score leaderboards of individual subtasks. In terms of the commonsense reasoning, the “ever-victorious generals” GPT-4V, WeMM, and XComposer-VL exceed other MLLMs, especially GPT-4V, which gets a score of 142.14. With respect to numerical calculation, GPT-4V still achieves first place, but falls short in the text translation. Regardless of whether it is commonsense reasoning, numerical calculation, or text translation, none of the highest scores exceed 150. This suggests that MLLMs have a lot of room for improvement in these capabilities. For the code reasoning, GPT-4V achieves a high score of 170, far ahead of other counterparts. For all of the cognition tasks, GPT-4V, Lion, and WeMM win the gold, silver, and bronze medals respectively, as shown in Fig. 2 (2).
@@ 4 Analysis
The third problem is a lack of reasoning. In the third row of Fig. 4, we can see from the red text that the MLLM already knows that the first image is not an office place, but still gives an incorrect answer of “yes”. Analogously, in the second image, the MLLM has calculated the right arithmetic result, but finally delivers a wrong answer. These phenomena indicate that the logic chain is broken during the reasoning process of MLLMs. Adding CoT prompts, such as “Let’s think step by step” [13], may yield better results. We look forward to a further in-depth research.
The fourth problem is object hallucination following instructions, which is exemplified in the fourth row of Fig. 4. When the instruction contains descriptions of an object that does not appear in the image, the MLLM will imagine that the object exists and ultimately gives a “yes” answer. Such a case of constantly answering “yes” results in an accuracy about 50% and an accuracy+ about 0, as shown in Tables 1 and 2. This suggests an urgent need to suppress hallucinations, and the community should take into account of the reliability of the generated answers.
@@ 5 Conclusion
This paper has presented the first MLLM evaluation benchmark MME that has four distinct characteristics in terms of task type, data source, instruction design, quantitative statistics. 30 advanced MLLMs are evaluated on MME and the experimental results show that there is still a large room to improve. We also summarize the common problem raised in experimental results, providing valuable guidance for the development of MLLM.